Cherry Corporation, a calendar year C corporation, is formed and begins business on 8/1/2019. In connection with its formation, Cherry incurs organizational expenditures of $53,800.

Round the per month amount to two decimal places. Round your final answer to the nearest dollar.

Determine Cherry Corporation's deduction for organizational expenditures for 2019. $______
Answer #1

1st year , least of following expenses considered as organizational expenses

Lesser of following:-

1) Actual Expenses incurred or

2) $5000, less Amount in Excess of $50000

Remaining balance Amount Amortised for 180 Months from Commencement of Business.

total Expenses= 53800

1) $53800

2) $5000 - (53800-50000)

5000 - 3800= $1200

Balance amount is Amortised for 180 Months:- 53800 - 1200= 52600

Business begins on August so August to December = 5 months

First Year allowed Expenses:-

1200 + (52600/180 X 5)

1200 + 1461= 2661